OpenStudy (anonymous):

What is the factored form of 2x2 − 7x + 6?

OpenStudy (anonymous):

hey is this the problem ? 2x^2-7x+6

OpenStudy (anonymous):

yes. (2x + 3)(x − 5) (2x − 3)(x − 2) (2x − 5)(x − 2) (2x + 1)(x − 2)

OpenStudy (anonymous):

these are my choices btw

OpenStudy (anonymous):

what do you think it is

OpenStudy (anonymous):

um 2x+3 x-5. ?

OpenStudy (anonymous):

Factorizing 2x^2-7x+6 = (2x-3)*(x-2)

OpenStudy (anonymous):

Factorizing x^4+2x^3-3x^2 Take out the common factor which in this case is x^2. Therefore it now = x^2(x^2+2x-3) You can also factorize the quadratic term x^2+2x-3 which = (x+3)*(x-1)

OpenStudy (anonymous):

yea I was thinking B

OpenStudy (anonymous):

yup! thank you!!

OpenStudy (anonymous):

so your answer would be_________

OpenStudy (anonymous):

2x-3 x-2

OpenStudy (anonymous):

sorry i got mixed up yes your correct